OpenStudy (anonymous):

Which is the cost per square inch of a pizza with diameter 20 inches if the cost of the pizza is $16? Be sure to round your answer to the nearest cent. about 1 cent about 5 cents about twenty-five cents about eighty cents confused help please

OpenStudy (phi):

Do you know how to find the area of a circle ?

OpenStudy (anonymous):

pi times radius squared

OpenStudy (phi):

ok, if the diameter is 20, what is the radius ?

OpenStudy (anonymous):

10?

OpenStudy (phi):

yes, radius is half of a diameter. now you need to find the area of a circle with r=10 A= pi*r*r = pi*10*10 pi is about 3.1416 (or use a calculator which uses more digits) A = 3.1416 *100 can you find the area ?

OpenStudy (anonymous):

314

OpenStudy (mathstudent55):

Correct. Now can you write $16 in cents?

OpenStudy (phi):

314.16 (you should keep more of the digits)

OpenStudy (anonymous):

0.16

OpenStudy (phi):

now you need cost per square inch that means cost divided by area can you do that ?

OpenStudy (mathstudent55):

No. There are 100 cents in $1, so in $16 dollars there are 16 * 100 cents.

OpenStudy (anonymous):

19.625

OpenStudy (phi):

can you show your work ?

OpenStudy (phi):

what is cost divided by area ? the cost is $16. the area is 314.16 square inches can you divide cost by area to get dollars per square inch ?

OpenStudy (anonymous):

19.635

OpenStudy (phi):

How. Can you show your work ?

OpenStudy (phi):

I think you are doing area divided by cost, which is "upside down"

OpenStudy (phi):

cost per square inch means cost / area NOT area/cost